# Spin-Orbit Misalignments of Eccentric Black Hole Mergers in AGN Disks

### v1

@@ 2.3 Spin-Orbit Tilt Evolution of Dynamical Mergers in the Disk
Fig. 1 shows a schematic representation of the evolution of the binary angular momentum for a BBH merger resulting from a three-body interaction embedded in an AGN disk. The left panel illustrates the initial configuration of the triple system: the BBH lies in the disk mid-plane (possibly having formed through gas-assisted dynamical formation), with the incoming single approaching at inclination $\varphi$ relative to the disk plane. Before the interaction, the binary angular momentum $\vec{J}_{\rm B}$ is assumed to be aligned with the overall direction of the disk angular momentum $\hat{J}_{\rm AGN}$. The individual BH spins are also assumed to be aligned, as accretion torques are expected to bring them into alignment with the AGN disk angular momentum vector on short timescales (e.g. Tagawa et al., 2020b; Vajpeyi et al., 2022; Santini et al., 2023; Mishra & Calcino, 2024). After the encounter (middle panel), the binary receives a dynamical kick, producing a misaligned spin–orbit tilt $\theta_{B}$ relative to $\hat{J}_{\rm AGN}$. While $\theta_{B}$ characterizes the orientation of the binary orbital angular momentum, its cosine directly enters the definition of the effective spin parameter $\chi_{\rm eff}$ under the assumption that individual black-hole spins are aligned with the AGN disk. Thus, the distribution of $\theta_{B}$ maps directly into the expected distribution of $\chi_{\rm eff}$. As the tilted binary evolves toward merger (right panel), it accretes gas, and some angular momentum $\Delta\vec{J}_{\rm gas}$ is added to $\vec{J}_{\rm B}$, driving the system back towards alignment.
@@ 2.4 Alignment prescription
We define $\dot{J}_{z}$ as the rate of change for the binary angular momentum component in the direction of the AGN disk angular momentum $\hat{J}_{\rm AGN}=\hat{J}_{z}$. Following the prescription of Tagawa et al. (2020b), we write
@@
where $a_{\rm bin}$ is the binary semi-major axis, $v_{\rm bin}$ its orbital velocity, $\dot{M}$ the mass accretion rate onto the binary, and $f_{\rm rot}$ a dimensionless parameter characterizing the typical angular momentum exchange required to accrete a parcel of gas. We treat $\dot{J}_{z}$ as constant during the alignment process, allowing us to define a characteristic alignment timescale in terms of the binary and disk properties, later expressed through Eq. (12). Similar to the method used in Fabj et al. (2025) to map the different regimes of stellar gas accretion in AGNs, we consider different regimes inside the disk to properly describe $\dot{M}$. We compare the size of the Bondi radius
@@
where $m_{\rm bin}$ is the mass of the binary, $c_{s}$ is the local sound speed in the disk, and $\Omega_{\bullet}$ is the binary orbital frequency around the SMBH. When $R_{B}<R_{H}$, we refer to the binary as being in an accretion dominated (AD) regime of the disk. On the other hand, we refer to the binary as being in a tidally limited (TL) regime when $R_{H}<R_{B}$. It follows that the mass accretion rate onto the binary can be approximated as
@@
where $\rho_{\rm g}$ is the disk density. The accretion rate is set by the smaller of the two characteristic radii, since gas accretion is limited by the more restrictive scale. In order to properly reflect that accretion begins at the boundary set by the minimum value between $R_{B}$ and $R_{H}$, we further restrict our analysis to binaries that are well embedded within their Hill sphere. This ensures we do not overestimate the accretion rate for binaries whose semi-major axis would otherwise extend close to the Hill radius. When $\dot{M}$ is in the accretion dominated regime we model it as $\dot{M}_{\rm AD}=0.8\,\dot{M}$, corresponding with recent hydrodynamical simulations of inclined (Dittmann et al., 2024) and eccentric (Calcino et al., 2024) binaries. However, we note that this generally corresponds to accretion rates far exceeding than the binary’s Eddington limit (see Section 5).
We adopt the parameter $f_{\rm rot}$, similar to the analysis of BH spin evolution in AGNs from Tagawa et al. (2020b), to parametrically quantify the magnitude of angular momentum transfer—from the accreting gas to the binary orbit—required to sustain accretion rate $\dot{M}$. For $f_{\rm rot}=1$ the transferred angular momentum corresponds to that of a circular orbit at $a_{\rm bin}$. A lower value of $f_{\rm rot}$ (e.g., 0.1) corresponds to less torque applied to the binary, for say, less rotationally supported gas. On the other hand, a very high efficiency (e.g., $f_{\rm rot}$ = 10), translates to ten times more angular momentum transport per unit mass than the circular Keplerian case. The hydrodynamical simulations in Dittmann et al. (2024) investigate the evolution of initially circular binaries placed at a small inclination (both highly prograde and retrograde) with respect to the disk mid-plane. By taking the ratio of the reported rate of inclination change to the initial tilt angle, we infer a characteristic alignment timescale and compare it with Eq. (1). We find that the Dittmann et al. (2024) results correspond to a value of $f_{\rm rot}=0.1$ in our model, which we identify as our fiducial case. However, since our prescription is phenomenological and neglects some physical processes (see Sec. 5),we also include results for $f_{\rm rot}=1,10$ to account for the possibility that the chosen prescription underestimates the alignment efficiency. In this sense, $f_{\rm rot}$ can be interpreted as a simple parameterization of the gas inflow into the binary, ranging from less efficient to more efficient alignment. Further hydrodynamical simulations will be necessary to determine which regime is the most representative of actual AGN disk conditions.
@@ 3.1 Gravitational Wave Timescale
Though we evaluate both the alignment and coalescence timescales starting from the binary properties at formation, throughout this paper we express our results in terms of the eccentricity when the GW signal enters the LVK band, defined at $f_{\rm gw}=10\,\rm Hz$. The GW peak frequency is the characteristic frequency where most of the gravitational wave power is radiated (D’Orazio & Samsing, 2018b). For three-body mergers, where the binary begins inspiraling while the third body is still bound, the systems are typically formed with very high eccentricities (e.g. Samsing, 2018; Samsing et al., 2022). In such cases, gravitational wave emission is strongly concentrated near pericenter, and the corresponding pericenter distance
@@ 4.1 Mapping Alignment Based on Disk Location
Figure 4: Ratio of negative to positive effective spin orientations, $N_{(\chi_{\rm eff}<0)}/N_{(\chi_{\rm eff}>0)}$, as a function of eccentricity at 10 Hz ($e_{\rm 10Hz}$). Results are shown for different alignment efficiencies: $f_{\rm rot}=0$ (no alignment, blue), $f_{\rm rot}=0.1$ (pink), and $f_{\rm rot}=1$ (green). For small eccentricities and high $f_{\rm rot}$, alignment is efficient and the ratio remains well below unity, reflecting the expected positive skewness of $\chi_{\rm eff}$ in AGN disks. For $f_{\rm rot}=0.1$ and at larger $e_{\rm 10Hz}$, the ratio grows and saturates near unity, indicating that alignment becomes less efficient than the GW inspiral. Without the effect of alignment, the excess of retrograde binaries reflects the outcome of three-body dynamics, while larger encounter angles ($\varphi$) suppress their production.
@@ 4.2 Spin-Orbit Tilt Distribution and Evolution
It is important to stress the assumptions implicit in this treatment. First, when evaluating Eq. (20) we adopt the binary separation $a_{0}$ at the moment of formation. This ignores the fact that gas can drive rapid orbital shrinkage (e.g. Baruteau et al., 2010; Li & Lai, 2023; Rowan et al., 2023, 2025b), which would reduce the merger time and shrink the window for alignment. Moreover, $e_{0}$ can also be modified by gas: if gas excites eccentricity in the binary, its merger time decreases even more steeply, further limiting alignment. On the contrary, if gas damps eccentricity, the inspiral timescale slows down, allowing alignment to proceed more efficiently. Simulations suggest that this behavior is determined by the orientation, where retrograde binaries exhibited eccentricity growth and vice versa for prograde binaries (e.g Rowan et al., 2023; Calcino et al., 2024; Dittmann et al., 2025). A more realistic calculation would couple the time evolution of $a_{0}$ and $e_{0}$ to the tilt evolution, solving the full system of differential equations rather than applying a single timescale, but at present this remains intractable. Second, as also observed in simulations (e.g. Dittmann et al., 2024) Eq. (20), like Eq. (13), assumes that retrograde binaries always evolve toward prograde orientations rather than becoming more retrograde, even for near-$180^{\circ}$ inclinations. The physical motivation for this choice is discussed further in Sec. 4.4.
@@
We also present results spanning the full range of eccentricities, focusing in particular on how the skewness of the effective spin parameter $\chi_{\rm eff}$, evolves depending on binary eccentricity at 10 Hz ($e_{\rm 10Hz}$) and on alignment efficiency $f_{\rm rot}$. By skewness we mean whether the distribution of $\chi_{\rm eff}$ is preferentially weighted toward positive (aligned) or negative (anti-aligned) values. As already discussed in Sec. 2.3, although the orbital inclination $\theta_{B}$ is not directly observable, if the gas in the disk aligns the individual BH spins with the disk angular momentum (e.g. Tagawa et al., 2020b), then the effective spin parameter becomes directly related to the binary tilt through $\chi_{\rm eff}\propto\cos\theta_{B}$ (e.g. Samsing et al., 2022). We therefore classify binaries with $\theta_{B}\leq 90^{\circ}$ as “positive” $\chi_{\rm eff}$ systems ($\chi_{\rm eff}>0$), and those with $\theta_{B}>90^{\circ}$ as “negative” $\chi_{\rm eff}$ systems ($\chi_{\rm eff}<0$).
@@
Finally, for higher alignment efficiency ($f_{\rm rot}=1$), the skewness of $\chi_{\rm eff}$ becomes strongly positive, recovering the distribution typically expected for AGN-assisted mergers where gas torques dominate (e.g. Tagawa et al., 2020b; McKernan et al., 2024; Ford & McKernan, 2025). Taken together, these results indicate that the interplay between binary eccentricity, disk density, and alignment efficiency $f_{\rm rot}$ controls whether the $\chi_{\rm eff}$ distribution in AGN disks is biased toward positive as generally expected or if its values remain closer to isotropic. The systems that merge before full alignment, are in addition expected to exhibit enhanced in-plane spin components $\chi_{P}$ (Ford & McKernan, 2025). In these cases, the effective spin can remain small even while the total spin magnitude is large, thus providing a complementary diagnostic of possible distinctive features of the AGN channel.
@@ 4.3 Upper Limit for the Alignment Angle
In this section we further explore how the efficiency of gas-driven alignment depends on the eccentricity at 10 Hz. Fig. 5 shows the initial distribution of binary orientations, $\theta_{B,0}$ (blue stars), compared to their final orientations after realignment $\theta_{B,\rm merge}$ (red stars). The black and red curves indicate the final alignment angle for a binary starting at $\theta_{B,0}=180^{\circ}$ (from Eq. 20), as a function of $e_{\rm 10Hz}$ for $f_{\rm rot}=1,10$. The shaded regions show the set of possible observed $\theta_{B,{\rm merge}}-e_{\rm 10Hz}$ combinations given the two values of $f_{\rm rot}$. Fig. 5 shows that for the highest value of $f_{\rm rot}$ binaries with $e_{\rm 10Hz}\lesssim 0.2$ are efficiently aligned regardless of their initial inclination, but for lower values of $f_{\rm rot}$ nearly the entire space remains possible. For both torque efficiencies the maximum post-alignment angle increases with eccentricity, further indicating that eccentric binaries are more likely to retain a spin-orbit misalignment at merger. In particular, the binaries with peak frequency above 10 Hz at formation also inspiral so rapidly that they can retain significant retrograde orientations despite strong alignment torques. The efficiency of this process, though, always depends sensitively on the local disk density: in denser regions alignment can occur much faster as $\tau_{\rm align}\propto\rho_{\rm g}$.
@@
Figure 5: Initial (blue stars) and final (red stars) inclination distributions $\theta_{B}$ as a function of $e_{\rm 10Hz}$, shown for $f_{\rm rot}=$ 1 (black), 10 (red). The black and red curves represent the maximum final inclination available for a binary starting at $\theta_{B,0}=180^{\circ}$, and the shaded region illustrates the available combinations of $\theta_{B,\,{\rm merge}}-e_{\rm 10Hz}$ at $f_{\rm rot}=10$. For the higher efficiency case, binaries with $e_{\rm 10Hz}\lesssim 0.2$ are fully realigned, while higher eccentricities lead to progressively larger residual tilts. The most eccentric mergers, with GW peak frequencies already above 10 Hz at formation, retain significant retrograde orientations due to their extremely short merger timescales.

### v2

@@ 2.3 Spin-Orbit Tilt Evolution of Dynamical Mergers in the Disk
Fig. 1 shows a schematic representation of the evolution of the binary angular momentum for a BBH merger resulting from a three-body interaction embedded in an AGN disk. The left panel illustrates the initial configuration of the triple system: the BBH lies in the disk mid-plane (possibly having formed through gas-assisted dynamical formation), with the incoming single approaching at inclination $\varphi$ relative to the disk plane. Before the interaction, the binary angular momentum $\vec{J}_{\rm B}$ is assumed to be aligned with the overall direction of the disk angular momentum $\hat{J}_{\rm AGN}$. The individual BH spins are also assumed to be aligned, as accretion torques are expected to bring them into alignment with the AGN disk angular momentum vector on short timescales (e.g. Tagawa et al., 2020a; Vajpeyi et al., 2022; Santini et al., 2023; Mishra and Calcino, 2024). After the encounter (middle panel), the binary receives a dynamical kick, producing a misaligned spin–orbit tilt $\theta_{B}$ relative to $\hat{J}_{\rm AGN}$. While $\theta_{B}$ characterizes the orientation of the binary orbital angular momentum, its cosine directly enters the definition of the effective spin parameter $\chi_{\rm eff}$ under the assumption that individual black-hole spins are aligned with the AGN disk  [^1] This assumption neglects possible misalignment of individual black-hole spins due to turbulent accretion via mini-disks or due to supersonic motion of the binary relative to the AGN disk, which can reorient mini-disks (e.g., Li et al., 2022a; Chen et al., 2022). . Thus, the distribution of $\theta_{B}$ maps directly into the expected distribution of $\chi_{\rm eff}$. As the tilted binary evolves toward merger (right panel), it accretes gas, and some angular momentum $\Delta\vec{J}_{\rm gas}$ is added to $\vec{J}_{\rm B}$, driving the system back towards alignment.
@@ 2.4 Alignment prescription
We define $\dot{J}_{z}$ as the rate of change for the binary angular momentum component in the direction of the AGN disk angular momentum $\hat{J}_{\rm AGN}=\hat{J}_{z}$. Following the prescription of Tagawa et al. (2020a), we write
@@
where $a_{\rm bin}$ is the binary semi-major axis, $v_{\rm bin}$ its orbital velocity, $\dot{M}$ the mass accretion rate defined later, and $f_{\rm rot}$ a dimensionless parameter characterizing the typical angular momentum exchange of a gas parcel. $\dot{J}_{z}$ is constant during the alignment process, allowing us to define a characteristic alignment timescale in terms of the binary and disk properties, later expressed through Eq. (12). Similar to the method used in Fabj et al. (2025) to map the different regimes of stellar gas accretion in AGNs, we set a control radius and consider different regimes inside the disk to properly describe $\dot{M}$. We compare the size of the Bondi radius
@@
where $m_{\rm bin}$ is the mass of the binary, $c_{s}$ is the local sound speed in the disk, and $\Omega_{\bullet}$ is the binary orbital frequency around the SMBH.
We note that using the total sound speed implicitly assumes optically thick, slow-diffusion conditions under which the radiation field remains coupled to the gas (Chen et al., 2024). In SG disks, radiation pressure is important mainly in the innermost regions (typically $\leq 10^{3}R_{g}$, e.g Goodman and Tan 2004; Gangardt et al. 2024), encompassing only a small fraction of embedded stellar-mass black holes (e.g. Stone et al., 2017) and where the optical depth is high. The fast-diffusion regime, in which radiation pressure decouples from the gas ($\tau\lesssim c/c_{s}$), is not modeled here, as it mainly occurs in the outer disk where gas pressure dominates and radiation pressure does not significantly influence the Bondi radius.
When $R_{B}<R_{H}$, we refer to the binary as being in an accretion-dominated (AD) regime of the disk. On the other hand, we refer to the binary as being in a tidally limited (TL) regime when $R_{H}<R_{B}$. It follows that the mass accretion rate onto the binary can be approximated as
@@
where $\rho_{\rm g}$ is the disk density. The accretion rate is set by the smaller of the two characteristic radii, since gas accretion is limited by the more restrictive scale. In order to properly reflect that accretion begins at the boundary set by the minimum value between $R_{B}$ and $R_{H}$, we further restrict our analysis to binaries that are well embedded within their Hill sphere. This ensures we do not overestimate the accretion rate for binaries whose semi-major axis would otherwise extend close to the Hill radius.
When $\dot{M}$ is in the accretion dominated regime we model it as $\dot{M}_{\rm AD}=0.08\,\dot{M}$, corresponding with recent hydrodynamical simulations of inclined (Dittmann et al., 2024) and eccentric (Calcino et al., 2024) binaries. However, we note that this generally corresponds to accretion rates far exceeding the binary’s Eddington limit (see Section 5), so we do not explicitly assume that the inflowing gas is deposited into the black holes. Instead, similar to the analysis of Tagawa et al. (2020a), we quantify the angular momentum exchange per inflowing gas parcel phenomenologically through the parameter $f_{\rm rot}$ where for $f_{\rm rot}=1$ the transferred angular momentum corresponds to that of a circular orbit at $a_{\rm bin}$. A lower value of $f_{\rm rot}$ (e.g., 0.1) corresponds to less torque applied to the binary, for say, less rotationally supported gas. On the other hand, a very high efficiency (e.g., $f_{\rm rot}$ = 10), translates to ten times more angular momentum transport per unit mass than the circular Keplerian case.
The hydrodynamical simulations in Dittmann et al. (2024) investigate the evolution of initially circular binaries placed at a small inclination (both highly prograde and retrograde) with respect to the disk mid-plane. By taking the ratio of the reported rate of inclination change to the initial tilt angle, we infer a characteristic alignment timescale and compare it with Eq. (1). We find that the Dittmann et al. (2024) results correspond to a value of $f_{\rm rot}=0.1$ in our model, which we identify as our fiducial case. However, since our prescription is phenomenological and neglects some physical processes (see Sec. 5),we also include results for $f_{\rm rot}=1,10$ to account for the possibility that the chosen prescription underestimates the alignment efficiency. In this sense, $f_{\rm rot}$ can be interpreted as a simple parameterization of the gas inflow into the binary, ranging from less efficient to more efficient alignment. Further hydrodynamical simulations will be necessary to determine which regime is the most representative of actual AGN disk conditions.
@@ 3.1 Gravitational Wave Timescale
Though we evaluate both the alignment and coalescence timescales starting from the binary properties at formation, throughout this paper we express our results in terms of the eccentricity when the GW signal enters the LVK band, defined at $f_{\rm gw}=10\,\rm Hz$. The GW peak frequency is the characteristic frequency where most of the gravitational wave power is radiated (D’Orazio and Samsing, 2018b). For three-body mergers, where the binary begins inspiraling while the third body is still bound, the systems are typically formed with very high eccentricities (e.g. Samsing, 2018; Samsing et al., 2022). In such cases, gravitational wave emission is strongly concentrated near pericenter, and the corresponding pericenter distance
@@ 4.1 Mapping Alignment Based on Disk Location
Figure 4: Ratio of negative to positive effective spin orientations, $N_{(\chi_{\rm eff}<0)}/N_{(\chi_{\rm eff}>0)}$, as a function of eccentricity at 10 Hz ($e_{\rm 10Hz}$). Results are shown for different alignment efficiencies: $f_{\rm rot}=0$ (no alignment, blue), $f_{\rm rot}=0.1$ (pink), and $f_{\rm rot}=1$ (green). For small eccentricities and high $f_{\rm rot}$, alignment is efficient and the ratio remains well below unity, reflecting a distribution weighted toward positive values of $\chi_{\rm eff}$ in AGN disks. For $f_{\rm rot}=0.1$ and at larger $e_{\rm 10Hz}$, the ratio grows and saturates near unity, indicating that alignment becomes less efficient than the GW inspiral. Without the effect of alignment, the excess of retrograde binaries reflects the outcome of three-body dynamics, while larger encounter angles ($\varphi$) suppress their production.
@@ 4.2 Spin-Orbit Tilt Distribution and Evolution
It is important to stress the assumptions implicit in this treatment. First, when evaluating Eq. (20) we adopt the binary separation $a_{0}$ at the moment of formation. This ignores the fact that gas can drive rapid orbital shrinkage (e.g. Baruteau et al., 2010; Li and Lai, 2023; Rowan et al., 2023, 2025b), which would reduce the merger time and shrink the window for alignment. Moreover, $e_{0}$ can also be modified by gas: if gas excites eccentricity in the binary, its merger time decreases even more steeply, further limiting alignment. On the contrary, if gas damps eccentricity, the inspiral timescale slows down, allowing alignment to proceed more efficiently. Simulations suggest that this behavior is determined by the orientation, where retrograde binaries exhibited eccentricity growth and vice versa for prograde binaries (e.g Rowan et al., 2023; Calcino et al., 2024; Dittmann et al., 2025). A more realistic calculation would couple the time evolution of $a_{0}$ and $e_{0}$ to the tilt evolution, solving the full system of differential equations rather than applying a single timescale, but at present this remains intractable. Second, as also observed in simulations (e.g. Dittmann et al., 2024) Eq. (20), like Eq. (13), assumes that retrograde binaries always evolve toward prograde orientations rather than becoming more retrograde, even for near-$180^{\circ}$ inclinations. The physical motivation for this choice is discussed further in Sec. 4.4.
@@
We also present results spanning the full range of eccentricities, focusing in particular on how the distribution of the effective spin parameter $\chi_{\rm eff}$, evolves depending on binary eccentricity at 10 Hz ($e_{\rm 10Hz}$) and on alignment efficiency $f_{\rm rot}$. We therefore examine whether the distribution of $\chi_{\rm eff}$ is preferentially weighted toward positive (aligned) or negative (anti-aligned) values. As already discussed in Sec. 2.3, although the orbital inclination $\theta_{B}$ is not directly observable, if the gas in the disk aligns the individual BH spins with the disk angular momentum (e.g. Tagawa et al., 2020a), then the effective spin parameter becomes directly related to the binary tilt through $\chi_{\rm eff}\propto\cos\theta_{B}$ (e.g. Samsing et al., 2022). We therefore classify binaries with $\theta_{B}\leq 90^{\circ}$ as “positive” $\chi_{\rm eff}$ systems ($\chi_{\rm eff}>0$), and those with $\theta_{B}>90^{\circ}$ as “negative” $\chi_{\rm eff}$ systems ($\chi_{\rm eff}<0$).
@@
Finally, for higher alignment efficiency ($f_{\rm rot}=1$), the distribution of $\chi_{\rm eff}$ values becomes strongly positive, recovering the one typically expected for AGN-assisted mergers where gas torques dominate (e.g. Tagawa et al., 2020a; McKernan et al., 2024; Ford and McKernan, 2025). Taken together, these results indicate that the interplay between binary eccentricity, disk density, and alignment efficiency $f_{\rm rot}$ controls whether the $\chi_{\rm eff}$ distribution in AGN disks is biased toward positive as generally expected or if its values remain closer to isotropic. The systems that merge before full alignment, are in addition expected to exhibit enhanced in-plane spin components $\chi_{P}$ (Ford and McKernan, 2025). In these cases, the effective spin can remain small even while the total spin magnitude is large, thus providing a complementary diagnostic of possible distinctive features of the AGN channel.
@@ 4.3 Upper Limit for the Alignment Angle
In this section we further explore how the efficiency of gas-driven alignment depends on the eccentricity at 10 Hz. Fig. 5 displays the initial distribution of binary orientations, $\theta_{B,0}$ (blue stars), and the corresponding final orientations at merger, $\theta_{B,\rm merge}$ (red stars), shown for the $f_{\rm rot}=10$ case. The red curve indicates the maximum final alignment angle for a binary starting at $\theta_{B,0}=180^{\circ}$ (from Eq. 20), as a function of $e_{\rm 10Hz}$, and the shaded red region illustrates the set of allowed $\theta_{B,{\rm merge}}-e_{\rm 10Hz}$ combinations for this torque efficiency. To illustrate how this behavior changes at lower torque efficiency, we also show the corresponding limiting curve and shaded region for $f_{\rm rot}=1$ in black.
@@
Fig. 5 shows that for the highest value of $f_{\rm rot}$ binaries with $e_{\rm 10Hz}\lesssim 0.2$ are efficiently aligned regardless of their initial inclination, but for lower values of $f_{\rm rot}$ nearly the entire space remains possible. For both torque efficiencies the maximum post-alignment angle increases with eccentricity, further indicating that eccentric binaries are more likely to retain a spin-orbit misalignment at merger. In particular, the binaries with peak frequency above 10 Hz at formation also inspiral so rapidly that they can retain significant retrograde orientations despite strong alignment torques. The efficiency of this process, though, always depends sensitively on the local disk density: in denser regions alignment can occur much faster as $\tau_{\rm align}\propto\rho_{\rm g}$.
The figure also includes GW190521, shown as a large green data-point with its associated uncertainties. The inclination is inferred from the reported $\chi_{\rm eff}$ in Abbott et al. (2020) assuming $\chi_{\rm eff}\propto\cos\theta_{B}$, while the eccentricity estimate is taken from Gayathri et al. (2022). Under these assumptions, GW190521 lies in a region of parameter space consistent with $f_{\rm rot}\sim 1$. More generally, the framework presented here can be applied to individual merger events by mapping constrained values of eccentricity and effective spin onto the predicted regions of the $e_{\rm 10Hz}-\chi_{\rm eff}$ plane. As illustrated in Fig. 5, even single-event measurements with large uncertainties can be used to assess whether a merger is consistent with rapid inspiral and incomplete alignment, or with more efficient alignment in the circular regime. We can therefore identify which values of the torque efficiency parameter $f_{\rm rot}$ are consistent with a given event. Since $f_{\rm rot}$ depends on local disk conditions, assuming a disk origin, we can also probe the local disk density at the merger site. Such applications will become increasingly powerful as eccentric waveform models improve and detector sensitivity increases.
@@
Figure 5: Initial (blue stars) and final (red stars) inclination distributions $\theta_{B}$ as a function of $e_{\rm 10Hz}$, with the final points shown for $f_{\rm rot}=10$. The red curve represents the maximum final inclination available for a binary starting at $\theta_{B,0}=180^{\circ}$, and the shaded region illustrates the available combinations of $\theta_{B,\,{\rm merge}}-e_{\rm 10Hz}$ at $f_{\rm rot}=10$. We also show the limiting curve for $f_{\rm rot}=$ 1 (black line), together with the associated black shaded region of available final combinations. For the higher alignment efficiency case, binaries with $e_{\rm 10Hz}\lesssim 0.2$ are fully realigned, while higher eccentricities lead to progressively larger residual tilts. The most eccentric mergers, with GW peak frequencies already above 10 Hz at formation, retain significant retrograde orientations due to their extremely short merger timescales. The green point shows GW190521. The inclination is inferred from the reported $\chi_{\rm eff}$, while the eccentricity estimate is taken from Gayathri et al. (2022)